# TSkips: Efficiency Through Explicit Temporal Delay Connections in Spiking Neural Networks

### v1

@@ Abstract
Spiking Neural Networks (SNNs) with their bio-inspired Leaky Integrate-and-Fire (LIF) neurons inherently capture temporal information. This makes them well-suited for sequential tasks like processing event-based data from Dynamic Vision Sensors (DVS) and event-based speech tasks. Harnessing the temporal capabilities of SNNs requires mitigating vanishing spikes during training, capturing spatio-temporal patterns and enhancing precise spike timing. To address these challenges, we propose TSkips, augmenting SNN architectures with forward and backward skip connections that incorporate explicit temporal delays. These connections capture long-term spatio-temporal dependencies and facilitate better spike flow over long sequences. The introduction of TSkips creates a vast search space of possible configurations, encompassing skip positions and time delay values. To efficiently navigate this search space, this work leverages training-free Neural Architecture Search (NAS) to identify optimal network structures and corresponding delays. We demonstrate the effectiveness of our approach on four event-based datasets: DSEC-flow for optical flow estimation, DVS128 Gesture for hand gesture recognition and Spiking Heidelberg Digits (SHD) and Spiking Speech Commands (SSC) for speech recognition. Our method achieves significant improvements across these datasets: up to 18% reduction in Average Endpoint Error (AEE) on DSEC-flow, 8% increase in classification accuracy on DVS128 Gesture, and up to  8% and  16% higher classification accuracy on SHD and SSC, respectively.
@@ 1 Introduction
Figure 1: Model size vs. average endpoint error (AEE) for fully spiking EV-FlowNet architectures [29] modified with Forward and Backward TSkips on the DSEC-flow dataset (lower AEE is better). Each tick on the horizontal axis represents a 3×3$\times$3 × increase in model size. TSkips achieve an average reduction of 15% in AEE and 40.75% in inference energy (E) while maintaining comparable AEE to larger baselines.
@@
Spiking Neural Networks (SNNs) are a class of bio-physically realistic models with Leaky Integrate-and-Fire (LIF) [1] neurons that offer a promising alternative for processing sequential data. By encoding information in the precise timing of spikes, SNNs capture spatio-temporal dependencies [43, 7] through the membrane potential of their LIF neurons. This intrinsic ability to process temporal information allows SNNs to excel at capturing dynamic patterns, effectively functioning as specialized RNNs [44] with reduced complexity.
@@
The ability of SNNs to leverage the precise timing of spikes offers a distinct advantage over traditional Artificial Neural Networks (ANNs) for sequential tasks. Note, ANNs are “stateless” in nature [15] and their reliance on specially curated input encoding schemes [59, 38] hinder their ability to accurately capture precise timing information crucial in event streams, as generated by Dynamic Vision Sensors (DVS) [33, 4]. Standard Recurrent Neural Networks (vRNNs) [47] and Long Short-Term Memory (LSTM) networks [19, 40] attempt to address this, but remain difficult to train [39] and computationally expensive. Furthermore, transformers and spiking transformers [31, 57] excel at sequential processing but come at the cost of significantly larger models and computational overhead.
@@
Deep SNNs offer the appealing combination of rich temporal information processing and computational efficiency, but training them effectively presents unique challenges [45]. The combination of multiple time steps with Back Propagation through Time (BPTT), non-differentiable spikes and sparse event data can worsen the vanishing spike problem [6]. To circumvent this, previous works have proposed adaptive LIF neurons with trainable threshold and leak [29] and approximate surrogate gradients [11, 35]. Other works [50, 52, 16] have considered synaptic delays as an additional parameter alongside weights, thus offering a richer representation of the spike patterns present in event data. We posit that these learnable synaptic delays only capture spatio-temporal patterns over a short window of time. Note, however, event data for speech recognition and optical flow also have long-term temporal dependencies that need to be captured through better spike representation.
@@
In contrast, this work introduces explicit temporal delays in forward and backward skip connections, named temporal skips (TSkips), for SNNs and hybrid ANN-SNN [36] models. TSkips offer finer control over spike timing, enhance responsiveness to temporal patterns, mitigate vanishing spikes, all while capturing long-term spatio-temporal dependencies for event-based data. In addition, our method utilizes adaptive LIF neurons [29] that have learnable leaks and thresholds that capture local temporal patterns.
@@
The next question to address would be, what are the optimal TSkip configurations within the network architecture and what are the corresponding temporal delays associated with them? To efficiently identify optimal TSkip configurations across different architectures and sequence lengths, we leverage training-free Neural Architecture Search (NAS) tailored for SNNs [27]. TSkips introduces minimal increase in model size with very few additional trainable parameters, minimal overhead and improves inference on event-based tasks. In addition, we have observed that training time is faster for the proposed networks compared to standard SNNs, RNNs and LSTMs. The contributions of this work can be summarized as follows:
@@
  • We analyzed various TSkip configurations, revealing the complexities of the search space, particularly the interplay between the temporal delay $(\Delta t)$, TSkip position and network depth. This complexity motivated the use of NAS to efficiently identify optimal TSkip architectures. (Section 4.2)
@@ 2.1 Training Deep SNNs
Early methods to train SNNs relied on ANN-to-SNN conversion [6, 37]. However, this approach often struggled to match ANN performance on complex sequential tasks [10], particularly with event data. A key obstacle in training deep SNNs is the non-differentiable nature of spike activations [11], which prevents the direct application of gradient-based optimization techniques, leading to subpar performance.
To overcome this, approximate surrogate gradients [11, 35] have been proposed, that replace non-differentiable spike activations with smooth gradient approximations, enabling the use of BPTT [53] to train deep SNNs. However, this approach introduces other challenges like vanishing gradients, especially with long sequences and sparse event data [6], and reduce accuracy compared to ANNs (due to approximate gradients). To address these challenges, previous work has investigated two main avenues: adaptive LIF neurons with learnable thresholds and leaks [29], and traditional skip connections (without delays) in SNNs [2]. While adaptive LIF neurons allow for dynamic adjustment and better capture of spatio-temporal patterns, and skip connections can improve accuracy and efficiency, neither approach fully addresses the challenges of representing and capturing complex temporal patterns. To achieve this, we propose TSkips, incorporating explicit temporal delays into the network architecture with forward and backward skip connections.
@@ 2.2 Delays in SNNs
The importance of precise timing information in event-based data, such as that from dynamic vision sensors, has motivated research into effectively capturing temporal dynamics in SNNs [43, 7]. Some works have introduced delays alongside the weights in SNNs to offer a richer representation of the spike patterns. For instance, [50] introduced a fixed delay in SNNs to improve learning of temporal complexities, [52] extended this by making these delays learnable and introducing a layer-specific maximum delay. [16] further refined this by making the per-layer maximum delay learnable and utilizing Dilated Convolutions with Learnable Spacings (DCLS) [25] to capture spatio-temporal patterns. However, learning these delays, as in [52, 16], introduces a new set of parameters with the same dimension as the weights, essentially doubling the number of trainable parameters. Additionally, these methods only capture local temporal patterns over a short window of time, due to their reliance on the dynamics of LIF neurons.
In contrast to these methods that rely on learning delays, our proposed method introduces explicit temporal delays in the form of forward and backward skip connections (TSkips) within the SNN architecture. The use of TSkips with adaptive LIF neurons effectively captures both local and long-term spatio-temporal patterns with minimal additional computational overhead.
@@ 3.1 Input Representation and Sensors
For efficient computation, SNNs require data in the form of discrete spikes, mirroring communication in biological neurons. DVS [33] cameras, output asynchronous spikes representing real-time changes in pixel intensity. This eliminates redundant data transmission and achieves higher temporal resolution than traditional frame-based cameras or rate coding methods [6]. These cameras trigger events when the log intensity change at a pixel crosses a threshold ($\theta$[12], represented as $||log(I_{t})-log(I_{t-1})||\geq\theta$. Events are encoded in the Address Event Representation (AER) format as tuples $(x,y,t,p)$ — representing pixel coordinates, timestamp, and polarity. Similarly, event-based audio datasets [8] mimic human auditory spiking activity, with spikes represented as $(x,t)$ — denoting spike unit and timestamp.
@@ 3.2 Neuron Model
The fundamental computational unit of SNNs is the spiking neuron. We employ Leaky Integrate-and-Fire (LIF) [1] neurons for their biological plausibility and computational efficiency. The LIF neuron integrates input spike information over time, accumulating it in its membrane potential. This potential gradually decays, allowing controlled forgetting of less relevant information. The dynamics of the LIF neuron can be described by
@@
where $U_{l}^{t}$ represents the membrane potential of the neurons in layer l at time step t, $\lambda$ is the leak factor, controlling the decay rate of the membrane potential. $W_{l}$ is the weight matrix for neurons in layer $l$ and $V_{l}^{th}$ is the voltage threshold of layer $l$.
@@
Figure 3: Illustration of TSkips in an SNN. (a) A 3-layer SNN with forward and backward TSkips, annotated with their origin and destination layers and the delay $(\Delta t)$. (b) Unrolled SNN with $T=4$ with a forward TSkip and $\Delta t=3$ compared with a standard skip connection (no delay). (c) Unrolled SNN with $T=4$ with a backward TSkip and $\Delta t=3$ compared with a standard backward connection ($\Delta t=1$[47]
@@ 3.3 Proposed Method
Efficiently training deep SNNs is often hindered by vanishing gradients [6], which impede accurate spike propagation across time steps. To address this, we propose a novel method that augments skip connections in SNNs by introducing explicit temporal delays ($\Delta t$) within the connections. We refer to these augmented skip connections, which can be forward, backward, or a combination of both, as TSkips. The duration of these explicit delays is constrained by $t<\Delta t<T$, where $t$ is the current time step and $T$ is the sequence length of the data. This constraint on $\Delta t$ ensures that the network is not receiving information from the future.
To illustrate the information flow in a neural network with TSkips, let us consider how they modify the output of a layer. Using $h_{l}^{t}$ to denote the input to a layer $l$ at time step $t$, a forward/backward TSkip from layer $l\pm k$ at time step $t-\Delta t$ can be represented as:
@@
where $f_{l}$ is the affine function of layer $l$, $\oplus$ represents either concatenation [22] or element-wise addition [17] and $W_{s}$ is a fixed shortcut path matrix [18] that randomly selects channels in $h_{l\pm k}^{(t-\Delta t)}$ to match the dimensions of $h_{l}^{t}$. Eq. 3 demonstrates that the output of layer $l$ at time step t is now influenced by the output of layer $l\pm k$ at time step $(t-\Delta t)$. Here, $k$ represents the difference in layer indices between connected layers. For example, if $l=4$ and $k=2$, a forward TSkip connects layers 2 and 4 with delay $\Delta t$.
Our method, incorporating TSkips, is illustrated in Fig. 3. Fig. 3(a) shows a simple 3-layer SNN architecture, illustrating both forward and backward TSkips between adjacent layers. Each connection is annotated with the origin and destination layers and $(\Delta t)$. Fig. 3(b) depicts the temporal unrolling of the network over $T=4$ time steps. It contrasts a standard skip connection $(\Delta t=0)$ and a forward TSkip with $\Delta t=3$. Fig. 3(c) similarly illustrates a backward TSkip with $\Delta t=3$ and contrasts it with a standard backward connection that has a delay of $\Delta t=1$, typically seen in vRNNs [47]. This comparison emphasizes that TSkips allow for more flexibility and can thus capture longer temporal dependencies compared to standard RNNs. This flexibility comes with minimal overhead, as TSkips are identity mappings that introduces very few additional parameters and no computational complexity to the network.
TSkips enhance the ability of SNNs and hybrid models to process sparse event streams by facilitating the propagation of temporally relevant information across non-adjacent layers. By capturing these long-term dependencies, TSkips can improve the network’s ability to learn complex temporal patterns. Furthermore, the explicit temporal delays in TSkips offer finer control over spike timing within the network. This finer control enhances the network’s responsiveness to temporal patterns by allowing it to selectively integrate information from different time steps. Additionally, by enabling the network to access relevant information from the past, TSkips can mitigate the issue of vanishing spikes. To further enhance temporal representation, we introduce a learnable scaling factor, $\alpha$, that controls the ratio of data at the current time-step ($t$) and data from the delayed time step ($t-\Delta t$) within the TSkip. This is incorporated into the skip as shown in Eq. 4:

### v2

@@ Abstract
Spiking Neural Networks (SNNs) with their bio-inspired Leaky Integrate-and-Fire (LIF) neurons inherently capture temporal information. This makes them well-suited for sequential tasks like processing event-based data from Dynamic Vision Sensors (DVS) and event-based speech tasks. Harnessing the temporal capabilities of SNNs requires mitigating vanishing spikes during training, capturing spatio-temporal patterns and enhancing precise spike timing. To address these challenges, we propose TSkips, augmenting SNN architectures with forward and backward skip connections that incorporate explicit temporal delays. These connections capture long-term spatio-temporal dependencies and facilitate better spike flow over long sequences. The introduction of TSkips creates a vast search space of possible configurations, encompassing skip positions and time delay values. To efficiently navigate this search space, this work leverages training-free Neural Architecture Search (NAS) to identify optimal network structures and corresponding delays. We demonstrate the effectiveness of our approach on four event-based datasets: DSEC-flow for optical flow estimation, DVS128 Gesture for hand gesture recognition and Spiking Heidelberg Digits (SHD) and Spiking Speech Commands (SSC) for speech recognition. Our method achieves significant improvements across these datasets: up to 18% reduction in Average Endpoint Error (AEE) on DSEC-flow, 8% increase in classification accuracy on DVS128 Gesture, and up to $\sim 8\%$ and $\sim 16\%$ higher classification accuracy on SHD and SSC, respectively.
@@ 1 Introduction
Spiking Neural Networks (SNNs) are a class of bio-physically realistic models with Leaky Integrate-and-Fire (LIF) (Abbott, 1999) neurons that offer a promising alternative for processing sequential data. By encoding information in the precise timing of spikes, SNNs capture spatio-temporal dependencies (Rathi & Roy, 2020; Chowdhury et al., 2021; Wu et al., 2017) through the membrane potential of their LIF neurons. This intrinsic ability to process temporal information allows SNNs to excel at capturing dynamic patterns, effectively functioning as specialized RNNs (Rathi & Roy, 2024; Deng et al., 2022) with reduced complexity.
Figure 1: Model size vs. average endpoint error (AEE) for fully spiking EV-FlowNet architectures (Kosta & Roy, 2023) modified with Forward and Backward TSkips on the DSEC-flow dataset (lower AEE is better). Each tick on the horizontal axis represents a 3×3$\times$3 × increase in model size. TSkips achieve an average reduction of 15% in AEE and 40.75% in inference energy (E) while maintaining comparable AEE to larger baselines.
@@
The ability of SNNs to leverage the precise timing of spikes offers a distinct advantage over traditional Artificial Neural Networks (ANNs) for sequential tasks. Note, ANNs are “stateless” in nature (Hagenaars et al., 2021b) and their reliance on specially curated input encoding schemes (Zhu et al., 2018b; Paredes-Vallés & De Croon, 2021) hinder their ability to accurately capture precise timing information crucial in event streams, as generated by Dynamic Vision Sensors (DVS) (Lichtsteiner et al., 2008; Brandli et al., 2014). Standard Recurrent Neural Networks (vRNNs) (Rumelhart et al., 1986) and Long Short-Term Memory (LSTM) networks (Hochreiter, 1997; Ponghiran & Roy, 2021; Gehrig et al., 2021b) attempt to address this, but remain difficult to train (Pascanu et al., 2013) and computationally expensive. Furthermore, transformers and spiking transformers (Li et al., 2022a; Zhou et al., 2023; Li et al., 2023) excel at sequential processing but come at the cost of significantly larger models and computational overhead.
@@
Deep SNNs offer the appealing combination of rich temporal information processing and computational efficiency, but training them effectively presents unique challenges (Pfeiffer & Pfeil, 2018; Rathi et al., 2021). The combination of multiple time steps with Back Propagation through Time (BPTT), non-differentiable spikes and sparse event data can worsen the vanishing spike problem (Hagenaars et al., 2021a; Neftci et al., 2019). To circumvent this, previous works have proposed adaptive LIF neurons with trainable threshold and leak (Fang et al., 2021b; Kosta & Roy, 2023; Wang & Li, 2023) and approximate surrogate gradients (Fang et al., 2021a; Neftci et al., 2019). Other works (Shrestha & Orchard, 2018; Sun et al., 2023b; Hammouamri et al., 2023) have considered synaptic delays as an additional parameter alongside weights, thus offering a richer representation of the spike patterns present in event data.
@@
In contrast, this work introduces explicit temporal delays in forward and backward skip connections, named temporal skips (TSkips), for SNNs and hybrid ANN-SNN (Kugele et al., 2021; Negi et al., 2024) models. TSkips offer finer control over spike timing, enhance responsiveness to temporal patterns, mitigate vanishing spikes, all while capturing long-term spatio-temporal dependencies for event-based data. In addition, our method utilizes adaptive LIF neurons (Yin et al., 2020; Hagenaars et al., 2021a; Kosta & Roy, 2023) that have learnable leaks and thresholds that capture local temporal patterns.
@@
The next question to address would be, what are the optimal TSkips configurations within the network architecture and what are the corresponding temporal delays associated with them? To efficiently identify optimal TSkips configurations across different architectures and sequence lengths, we leverage training-free Neural Architecture Search (NAS) tailored for SNNs (Kim et al., 2022). TSkips introduces minimal increase in model size with very few additional trainable parameters, minimal overhead and improves inference on event-based tasks. In addition, we have observed that training time is faster for the proposed networks compared to standard SNNs, RNNs and LSTMs. The contributions of this work can be summarized as follows:
@@
  • We analyzed various TSkips configurations, revealing the complexities of the search space, particularly the interplay between the temporal delay $(\Delta t)$, TSkips position and network depth. This complexity motivated the use of NAS to efficiently identify optimal TSkips architectures. (Section 4.2)
@@ 2.1 Training Deep SNNs
Early methods to train SNNs relied on ANN-to-SNN conversion (Cao et al., 2015; Panda et al., 2020). However, this approach often struggled to match ANN performance on complex sequential tasks (Deng et al., 2020), particularly with event data. To overcome the challenges associated with ANN-to-SNN conversion, Kugele et al. (2020) focus on converting existing neural network architectures into spiking neural networks using streaming rollouts (Fischer et al., 2018). By temporally unfolding the recurrent structures in the spiking domain the DenseNet skip connections essentially have a single step delay, much like vRNNs.
A key obstacle in training deep SNNs is the non-differentiable nature of spike activations (Fang et al., 2021a), which prevents the direct application of gradient-based optimization techniques, leading to subpar performance. To overcome this, approximate surrogate gradients (Fang et al., 2021a; Neftci et al., 2019; Wu et al., 2017) have been proposed, that replace non-differentiable spike activations with smooth gradient approximations, enabling the use of BPTT (Werbos, 1990) to train deep SNNs. However, this approach introduces other challenges like vanishing gradients, especially with long sequences and sparse event data (Hagenaars et al., 2021a; Neftci et al., 2019), and reduce accuracy compared to ANNs (due to approximate gradients). To address these challenges, previous work has investigated two main avenues: adaptive LIF neurons with learnable thresholds and leaks (Fang et al., 2021b; Kosta & Roy, 2023; Wang & Li, 2023), and skip connections in SNNs (Benmeziane et al., 2023). While adaptive LIF neurons allow for dynamic adjustment and better capture of spatio-temporal patterns, skip connections can improve accuracy and efficiency, However, neither approach fully addresses the challenges of representing and capturing complex temporal patterns. To achieve this, we propose TSkips, incorporating explicit temporal delays into the network architecture with forward and backward skip connections.
@@ 2.2 Delays in SNNs
The importance of precise timing information in event-based data, such as that from dynamic vision sensors, has motivated research into effectively capturing temporal dynamics in SNNs (Rathi & Roy, 2020; Chowdhury et al., 2021). SCTT (Soo et al., 2023) explore temporal skips in recurrent networks and are evalauted on cognitive tasks that have long-term temporal dependencies. However, SCTT employs an expensive evaluation strategy to identify optimal delay and temporal skip hyperparameters that involves training several models across multiple tasks to identify optimal connections. TTFS (Kim et al., 2024) explore adding delays in skip connections by encoding the latency of the first spike emitted by a neuron. This encoded delay is then introduced into skip connections in SNNs as learnable parameters that promote local path synchronization. However, the evaluation of both these works are limited to simple tasks that are not typically conducive to SNNs. Furthermore, while the spike encoding strategy of TTFS allows for better synchronization of spikes for the model, it does not account for factors such as noise or irregularities in the event stream or long periods of sparse events.
Other works have introduced delays alongside the weights in SNNs to offer a richer representation of the spike patterns. For instance, Shrestha & Orchard (2018) introduced a fixed delay in SNNs to improve learning of temporal complexities, Sun et al. (2023b) extended this by making these delays learnable and introducing a layer-specific maximum delay. Hammouamri et al. (2023) further refined this by making the per-layer maximum delay learnable and utilizing Dilated Convolutions with Learnable Spacings (DCLS) (Khalfaoui-Hassani et al., 2023) to capture spatio-temporal patterns. However, learning these delays, as in Sun et al. (2023b); Hammouamri et al. (2023), introduces a new set of parameters with the same dimension as the weights, essentially doubling the number of trainable parameters.
In contrast to these methods that rely on expensive tempral skip configuration evaluation strategies, estimating delay values for local path synchronization or learning delays, our proposed method introduces explicit temporal delays in SNNs. We explore these delays in both forward and backward temporal skip connections, named TSkips. The use of TSkips coupled with a learnable scaling factor and adaptive LIF neurons effectively captures long-term spatio-temporal patterns with minimal additional computational overhead.
@@ 3.1 Input Representation and Sensors
For efficient computation, SNNs require data in the form of discrete spikes, mirroring communication in biological neurons. DVS (Lichtsteiner et al., 2008) cameras output asynchronous spikes representing real-time changes in pixel intensity $(I_{t})$. This eliminates redundant data transmission and achieves higher temporal resolution than traditional frame-based cameras or rate coding methods (Cao et al., 2015). These cameras trigger events when the log intensity change at a pixel crosses a threshold ($\theta$(Gallego et al., 2022), represented as $||log(I_{t})-log(I_{t-1})||\geq\theta$. Events are encoded in the Address Event Representation (AER) format as tuples $(x,y,t,p)$ — representing pixel coordinates, timestamp, and polarity. Similarly, event-based audio datasets (Cramer et al., 2022) mimic human auditory spiking activity, with spikes represented as $(x,t)$ — denoting spike unit and timestamp.
@@ 3.2 Neuron Model
The fundamental computational unit of SNNs is the spiking neuron. We employ Leaky Integrate-and-Fire (LIF) (Abbott, 1999) neurons for their biological plausibility and computational efficiency. The LIF neuron integrates input spike information over time, accumulating it in its membrane potential. This potential gradually decays, allowing controlled forgetting of less relevant information. The dynamics of the LIF neuron can be described by
@@
where $U_{l}^{t}$ represents the membrane potential of the neurons in layer l at time step t, $\lambda$ is the leak factor, controlling the decay rate of the membrane potential. $W_{l}$ is the weight matrix for neurons in layer $l$, $V_{l}^{th}$ is the voltage threshold of layer $l$ and $O_{l}^{t}$ is the output generated by layer $l$ at time step t.
@@ 3.3 Proposed Method
Figure 3: Illustration of TSkips in an SNN. (a) A 3-layer SNN with forward and backward TSkips, annotated with their origin and destination layers and the delay $(\Delta t)$. (b) Unrolled SNN with $T=4$ with a forward TSkips and $\Delta t=3$ compared with a standard skip connection (no delay). (c) Unrolled SNN with $T=4$ with a backward TSkips and $\Delta t=3$ compared with a standard backward connection ($\Delta t=1$(Rumelhart et al., 1986)
Efficiently training deep SNNs is often hindered by vanishing gradients (Hagenaars et al., 2021a; Neftci et al., 2019), which impede accurate spike propagation across time steps. To address this, we propose a novel method that augments skip connections in SNNs by introducing explicit temporal delays ($\Delta t$) within the connections. We refer to these augmented skip connections, which can be forward, backward, or a combination of both, as TSkips. The duration of these explicit delays is constrained by $0<t-\Delta t<T$, where $t$ is the current time step and $T$ is the sequence length of the data. Specifically, $t$ represents the set of all discrete time steps within the sequence, denoted as $t=\left\{t_{i},\forall i=0,\cdots,T\right\}$. This constraint on $\Delta t$ ensures that the network does not receive future information.
To illustrate the information flow in a neural network with TSkips, let us consider how they modify the output of a layer. If $h_{l}^{t}$ denotes the input to a layer $l$ at time step $t$, then a forward TSkips provides input $h_{l-k}^{t}$ and a backward TSkips provides input $h_{l+k}^{t}$. For brevity, we denote these skip inputs as $h_{l\pm k}^{t}$, where $k$ indicates the layer separation between layer $l$ and the source of the skip connection. TSkips can be represented as:
@@
where $f_{l}$ is the affine function of layer $l$, $\oplus$ represents either concatenation (Huang et al., 2017) or element-wise addition (He et al., 2015) and $W_{s}$ is a fixed shortcut path matrix (He et al., 2016) that randomly selects channels in $h_{l\pm k}^{(t-\Delta t)}$ to match the dimensions of $h_{l}^{t}$. Eq. 3 demonstrates that the output of layer $l$ at time step $t$ is now influenced by the output of layer $l\pm k$ at time step $(t-\Delta t)$. Here, $k$ represents the difference in layer indices between connected layers. For example, if $l=4$ and $k=2$, a forward TSkips connects layers 2 and 4 with delay $\Delta t$.
Our method, incorporating TSkips, is illustrated in Fig. 3. Fig. 3(a) shows a simple 3-layer SNN architecture, illustrating both forward and backward TSkips between adjacent layers. Each connection is annotated with the origin and destination layers and $(\Delta t)$. Fig. 3(b) depicts the temporal unrolling of the network over $T=4$ time steps. It contrasts a standard skip connection $(\Delta t=0)$ and a forward TSkips with $\Delta t=3$. Fig. 3(c) similarly illustrates a backward TSkips with $\Delta t=3$ and contrasts it with a standard backward connection that has a delay of $\Delta t=1$, typically seen in vRNNs (Rumelhart et al., 1986). This comparison emphasizes that TSkips allow for more flexibility and can thus capture longer temporal dependencies compared to standard RNNs. This flexibility comes with minimal overhead, as TSkips are identity mappings that introduces very few additional parameters and no computational complexity to the network.
TSkips enhance the ability of SNNs and hybrid models to process sparse event streams by facilitating the propagation of temporally relevant information across non-adjacent layers. By capturing these long-term dependencies, TSkips can improve the network’s ability to learn complex temporal patterns. Furthermore, the explicit temporal delays in TSkips offer finer control over spike timing within the network. This finer control enhances the network’s responsiveness to temporal patterns by allowing it to selectively integrate information from different time steps. Additionally, by enabling the network to access relevant information from the past, TSkips can mitigate the issue of vanishing spikes. To further enhance temporal representation, we introduce a learnable scaling factor, $\alpha$, that controls the ratio of data at the current time-step ($t$) and data from the delayed time step ($t-\Delta t$) within the TSkips. This is incorporated into the skip as shown in Eq. 4: